OpenStudy (anonymous):

f(x) = e^2x / x^5 + 1 f ' x = ? This one seems reallly confusing

OpenStudy (accessdenied):

They can look quite confusing at first, but once you're comfortable with all the rules, they really aren't too bad. :) In this equation, we can see that the second term on the right-hand side is just going to fall off (derivative of a constant being 0), so we can ignore it. The first term is a quotient of two functions, so we may use the quotient rule, or if we change the exponent of the denominator to a negative and bring it back up, it becomes easier for us to use the product rule. \[e ^{2x}/x ^{5} = e^{2x}x^{-5}\] The product rule applied to this will be... \[(e^{2x} \times x^{-5})' = (e^{2x})'x^{-5} + e^{2x}(x^{-5})'\] From the property of e^x and its derivatives being equal, we just use the chain rule. So, the derivative of e^(2x) is 2e^(2x). From there, it is mostly just simplification. \[f'(x) = 2e^{2x}x^{-5} - 5e^{2x}x^{-6}\] Of course, you may simplify further (return the exponents to positives), but this should be sufficient. :)

OpenStudy (anonymous):

the answer should be in the form df/dx however that is

OpenStudy (accessdenied):

The f'(x) and df/dx mean the same thing. You can represent the derivative with either notation, so you just replace the "f'(x)" in the above equation with "df/dx".

OpenStudy (anonymous):

how do you simplify that, the homework is being stuborn

OpenStudy (accessdenied):

Well, because you have negative exponents, you would place them back in the denominator with positive exponents (just a property of exponents). \[df/dx = (2e^{2x}/x^{5}) - (5e^{2x}/x^{6})\] (Sorry, these should be fractions, but I don't know how to add them with the equation thing here) If you need to have one denominator, you just multiply the first term by (x/x) \[df/dx = (2xe^{2x} - 5e^{2x})/x^{6}, = [(2x - 5)e^{2x}]/x^{6}\] That should be completely simplified.

OpenStudy (anonymous):

i hate my homework